# DeltaSpace: A Semantic-aligned Feature Space for Flexible Text-guided Image Editing

### v2

@@ Abstract
Index Terms:
Text-guided image editing, generative adversarial networks, diffusion models.
@@ 2.3 Text-guided Image Editing with Diffusion Models
In this work, we propose a novel diffusion model, called style-conditioned diffusion model, which utilizing the semantically rich and comprehensive representation offered by StyleGAN as the conditional signal to control the model.
@@ 3.1.1 Discussion with the Original CLIP Space
Quantitative experiments in Tab. I also demonstrate that the straightforward way cannot ensure accurate editing. In contrast, our proposed DeltaSpace allows for accurate and flexible image editing, as shown in Fig. 5, demonstrating again DeltaSpace is more aligned than the original CLIP space.
@@ 4.2.1 Preliminaries
$x_{t-1}$ $=\sqrt{\alpha_{t-1}}\underbrace{\left(\frac{x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}\left(x_{t},t\right)}{\sqrt{\alpha_{t}}}\right)}_{\text{``predicted }x_{0}\text{ " }}$ (13)
@@ 5.2.2 Qualitative Comparison
For a fair comparison, we compare DeltaEdit-G with state-of-the-art GAN-based image editing methods, TediGAN [5], StyleCLIP-LM [7], StyleCLIP-GD [7] and StyleMC [9]. As shown in Fig. 11, the results produced by TediGAN are highly entangled and almost fail in editing attributes, such as “long hair” and “eyeglasses”. The results of StyleCLIP-LM and StyleMC are unstable and also entangled. Moreover, the results of StyleCLIP-GD have better disentanglement than StyleCLIP-LM as it is well-tunned for each case. However, the disadvantage of StyleCLIP-GD is the complex hyper-parameters tunning and long inference time as reported in Tab. V. In comparison, DeltaEdit-G yields the most impressive and disentangled results almost in all cases.
@@ 5.2.3 Quantitative Comparison
TABLE I: Quantitative comparison results of GAN-based methods.
@@
StyleCLIP-GD $\rho=2$ $12.06$ $22.31$ $0.86$
StyleCLIP-GD $\rho=3$ $16.85$ $19.50$ $0.77$
StyleCLIP-GD $\rho=4$ $22.85$ $17.61$ $0.68$
@@
DeltaEdit-G (ours) ($\mathcal{S}$ space) $18.88$ $13.49$ $0.61$
DeltaEdit-G (ours) ($\mathcal{W+}$ space) $21.63$ $12.83$ $0.55$
DeltaEdit-G (ours) (straightforward) $44.62$ $11.21$ $0.29$
In Tab. I, we present objective measurements of FID, PSNR, and IDS (identity similarity before and after editing by Arcface [66]) for GAN-based methods comparison. All results are the average on ten given texts. Compared with the state-of-the-art GAN-based methods, our DeltaEdit-G achieves the best performance on all metrics.
We also compare DeltaEdit-G with the straightforward solution to text-free training in Tab. I. For fairness, the straightforward solution is also trained in the $\mathcal{S}$ space and supervised by the same loss functions as the final solution. Compared with the straightforward solution, the final solution (ours) can largely improve the inference performance on all metrics, benefiting from the well-aligned CLIP Delta image-text space.
TABLE II: User preference study on editing accuracy (Acc.) and visual realism (Real.)
@@
Finally, we conduct human subjective evaluations upon the editing accuracy (Acc) and visual realism (Real). We compare DeltaEdit-G with the strong GAN-based baseline, StyleCLIP-GD, under different hyper-parameter $\rho$ settings. In total, 20 evaluation rounds are performed and 40 participants are invited. At each round, we present results of randomly sampled editing text to each participant. Participants were asked to choose the best text-guided image editing output considering Acc and to select the output images (not limited to 1 image) which are visually realistic (Real). The results are listed in Tab. II, showing the superiority of our model.
@@ 5.3.1 Qualitative Results
TABLE III: Quantitative results of image reconstruction using SSIM, LPIPS, and MSE metrics.
@@
Real Image Reconstruction. The quality of image reconstruction can reflect the accuracy of a generative model in encoding-decoding real images. As shown in Tab. III, we evaluate image reconstruction accuracy of various methods, including: (1) GAN inversion methods (including pretrained StyleGAN with $\mathcal{W}$ or $\mathcal{W+}$ sapce [16] as well as VQ-GAN [69]); (2) VAE-based methods (such as VQ-VAE2 [70]); (3) diffusion-based methods (including DDIM [52], DiffAE [53], and our proposed diffusion model). All models are trained on the FFHQ [15] dataset and tested on the CelebA-HQ [59] 30K dataset. All diffusion-based methods use T=100 for decoding. Our diffusion method and the diffusion-based DiffAE approach demonstrate superior performance, surpassing previous GAN-based and VAE-based methods by a large margin, thus confirming their near-perfect reconstruction capabilities.
@@
Figure 15: Style mixing results using our DeltaEdit-D approach. The first column shows the style images, while the first row displays the content images.
Style Mixing. We also evaluate our DeltaEdit-D on style mixing task [71], which aims to transfer the stylistic characteristics of a style image onto a content image. To achieve this, we apply the e4e encoder [57] to convert both the style image and the content image into corresponding latent codes. Subsequently, we replace the high-level latent code of the content image with the corresponding one from the style image. Finally, the mixed latent code is utilized to control the diffusion model to generate a novel image, as depicted in Fig. 15. The results show that the painting style of the style image (first column) is effectively blended while retaining most of the details of the content image (first row). This implies that the high-level GAN latent code can control the high-level semantics such as background, color, and lighting in the results generated by the diffusion model, while the low-level GAN latent code can effectively control the structure of the image, including pose, gender, and facial shape.
@@ 5.3.2 Qualitative Comparison
TABLE IV: Quantitative comparison results of our method with diffusion-based methods.
@@ 5.3.3 Quantitative Comparison
In Tab. IV, we present objective measurements of FID, PSNR, and IDS (identity similarity before and after editing by Arcface [66]) for diffusion-based methods comparison. The results are the average on ten given texts. The quantitative results have the best performance across multiple metrics. These results are consistent with the qualitative results that our DeltaEdit-D method exhibits the best disentanglement and identity preservation compared to diffusion-based models.
@@ 5.4 Comparison between DeltaEdit-G and DeltaEdit-D
5.4.1 Qualitative Comparison
5.4.2 Quantitative Comparison
In Tab. IV, we additionally present a quantitative comparison between DeltaEdit-G and DeltaEdit-D. The experiments are conducted by averaging results over the same ten text prompts. As shown in the table, DeltaEdit-D achieves better quality effects and higher identity similarity compared to DeltaEdit-G. This aligns with the qualitative results, demonstrating that by instantiated on the diffusion model, our DeltaEdit method can generate higher-quality edited outputs.
5.5 Model Analysis
5.5.1 Text-prompt Setting
TABLE V: Time efficiency comparison of our method and other state-of-the-art methods. * means that the additional time for manually hyper-parameters tuning during the inference is excluded, which is typically 8-9 seconds for each case.
@@ 5.5.1 Text-prompt Setting
5.5.2 Image Condition Setting
Considering that in the training phase $\Delta S$ is derived from images, we conduct a series of experiments to validate the effectiveness of image differences as editing conditions during testing: (1) Under conditions identical to training, we employ the same image as the source image and the input image to verify whether the proposed method can reconstruct the target image using image feature differences as conditions during testing. As shown in Fig. 20 (a), the editing results effectively learn the semantic direction from source to target images while achieving high-quality reconstruction, thereby validating the consistency between training and testing phases. Furthermore, Appendix B.3 demonstrates our method’s reconstruction errors during testing, where low latent-space MSE (2.32) and pixel-space MSE (0.17) quantitatively confirm the effectiveness of the approach. (2) Furthermore, we utilize distinct images as source, target, and input images. Fig. 20 (b) illustrates how our method successfully modifies the input image along the semantic direction of image differences while preserving identity-related information of the input image. (3) As shown in Fig. 20 (c), when strategically using image pairs differing in only one attribute, our method achieves precise target semantic editing comparable to text-guided single-attribute manipulation.
Figure 20: Editing results using different image pairs as editing conditions.
@@ 5.5.3 Efficiency Analysis
To validate the efficiency and flexibility of the proposed method, we compare the computation time with TediGAN, StyleMC and three StyleCLIP approaches in Tab. V. Specifically, TediGAN first requires 12+ hours to encode images and texts into a common space and then trains the encoding module. StyleCLIP-OP manipulates images with an optimization process, which requires several tens of seconds. StyleCLIP-LM is fast during inference, but needs 10-12 hours to train a mapper network for per text prompt. StyleCLIP-GD requires about 4 hours to pre-compute the global edit directions. However, since the pre-computed directions are rough and cannot be directly applied to control the manipulation, it requires additional manual tuning for different text prompts, which typically takes 8-9 seconds for each case. Differently, StyleMC is fast during training. However, for N text prompts, StyleMC still needs N times longer to train the corresponding model. On the contrary, our DeltaEdit-G method avoids the labor time during training and inference. Once trained, our Delta Mapper is universal and can directly manipulate images with new text prompts efficiently. Meanwhile, our Delta Mapper only needs 2-3 hours for training, since the training is directly conducted on the latent space without generating images in each iteration. However, it is worth noting that DeltaEdit-D requires additional training time to learn the style-based diffusion model, resulting in a longer training phase compared to DeltaEdit-G. Additionally, since DeltaEdit-D relies on iterative denoising via the diffusion model to generate edited results, its inference time is also longer. Nevertheless, this progressive denoising process enables DeltaEdit-D to achieve higher-quality image edits than DeltaEdit-G, as quantitatively demonstrated in Fig. 17 and Tab. IV.
@@ 5.5.4 Choice of the Editing Space
To find an appropriate space for the proposed method, we conduct experiments by performing editing in $\mathcal{W+}$ space and $\mathcal{S}$ space, validating the results on both GAN and diffusion models.. The quantitative and qualitative results in Tab. I and Fig. 19 show that $\mathcal{S}$ results can achieve better visual quality and identity preservation than $\mathcal{W+}$. For instance, in Fig. 19, when considering the attribute “bangs”, $\mathcal{W+}$ results succeed in adding bangs to given faces. However, the attribute-irrelevant contents are heavily changed, such as poses, skin color, and background. In contrast, $\mathcal{S}$ results have more disentanglement than $\mathcal{W+}$, e.g., the poses, facial identities are well preserved. Thus, we implement our DeltaEdit in $\mathcal{S}$ space, and introduce $R_{s}$ in the disentanglement part of Sec. 3.2 to further improve the disentanglement performance.
@@ 5.5.5 Effectiveness of Relevance Matrix $R_{s}$
For further improving the disentanglement, we introduce relevance matrix $R_{s}$ to limit some irrelevant channels from changing. The fourth row in Fig. 19 shows that, with $R_{s}$, our method can successfully edit the desired attributes while preserving the text-irrelevant content unchanged. For example, for “bangs”, the generated results have added bangs while accurately preserving the background and facial pose. Additionally, we conduct experiments on the straightforward solution of our proposed method to validate the effectiveness of $R_{s}$ in Fig. 21. As shown in the second column of Fig. 21, the straightforward method suffers from the “average face” problem. When $R_{s}$ is introduced (third column), this issue is largely alleviated. This improvement stems from Rs’s ability to constrain irrelevant values, which further verifies its effectiveness.
Figure 21: Effectiveness Experiments of $R_{s}$ with the input text “face with yellow hair”: (a) baseline( straightforward version), (b) straightforward version with $R_{s}$, and (c) final version. $R_{s}$ can further improve disentanglement by suppressing irrelevant attributes.
6 Limitations and Discussions
While our DeltaSpace framework demonstrates strong performance in semantic-driven editing, several limitations stem from its reliance on the CLIP embedding space. The semantic expressiveness of DeltaSpace is inherently constrained by CLIP’s training data distribution, particularly for rare or underrepresented attributes such as “blue hair” or “blue eyes”, where the model struggles to generate precise edits due to insufficient representation in the original embedding space. Furthermore, the framework exhibits degraded performance when applied to domains significantly different from CLIP’s training distribution or when handling novel attribute combinations absent in the training data. These limitations suggest opportunities for future improvements through domain-adaptive fine-tuning or hybrid modeling approaches that augment CLIP’s pretrained representations with more targeted semantic supervision.
Additionally, our method shares the same scalability limitation as other category-specific editing methods, requiring per-category training due to fundamental differences in StyleGAN’s style spaces across domains. This prevents cross-category generalization. Future work will investigate adapter networks to enable generalization across semantically similar categories while maintaining precision.
@@ 7 Conclusion
In this paper, we introduce a flexible text-guided image editing framework named DeltaEdit. This framework is built upon the well-aligned CLIP DeltaSpace we have identified. During the training phase, it learns to map CLIP visual feature differences to the latent space directions of a generative model. Subsequently, during the inference phase, the framework utilizes CLIP text feature differences to predict the latent space directions. In this way, it is trained in a text-free manner and eliminates the need for an expensive collection of image-text pairs. Once trained, it can be well generalized to any unseen text prompt for zero-shot inference. Extensive qualitative and quantitative experiments of DeltaEdit with different generative models, including the GAN model and the diffusion model, demonstrate the superiority of our method, in terms of the capability of generating high-quality results, the efficiency in the training and inference phase, and the generalization to any unseen text prompt.
8 Acknowledgements
The authors would like to thank Yujun Shen at Ant Group for his helpful discussions and suggestions on this work.

### v3

@@ Abstract
keywords:
Text-guided image editing, CLIP Space, generative adversarial networks, diffusion models
@@ 1 Introduction
The preliminary result of this work is published at [21]. Compared to the conference paper, we include the following new contents: (1) We provide a detailed analysis of DeltaSpace in Sec. 3.1. Based on the semantic-aligned properties of DeltaSpace, our DeltaEdit framework establishes effective, generalized and versatile editing performance on two types of generative models. The previous approach is only effective for the GAN model. (2) We propose a style-conditioned diffusion model in Sec. 4.2.2, which leverages the semantic space of StyleGAN to control forward and reverse processes of the conditional diffusion model. This combination achieves strong detailed reconstruction while yielding remarkable image editing, as shown in Fig. 2. (3) We include evaluations and comparisons of DeltaEdit-D to further validate the effectiveness and generalizability of the method in Sec. 5.3 and Sec. 5.4, including semantically meaningful latent interpolation, real image reconstruction, flexible text-guided editing, etc.
@@ 2.3 Text-guided Image Editing with Diffusion Models
In this work, we propose a novel diffusion model, named style-conditioned diffusion model, which utilizes the semantically rich and comprehensive representation offered by StyleGAN as the conditional signal to control the model.
@@ 3.1.1 Discussion with the Original CLIP Space
Quantitative experiments in Tab. 1 also demonstrate that the straightforward way cannot ensure accurate editing. In contrast, our proposed DeltaSpace allows for accurate and flexible image editing, as shown in Fig. 5, demonstrating again DeltaSpace is more aligned than the original CLIP space.
@@ 4.2.1 Preliminaries
[…]
@@ 5.2.2 Qualitative Comparison
For a fair comparison, we compare DeltaEdit-G with state-of-the-art GAN-based image editing methods, TediGAN [1], StyleCLIP-LM [9], StyleCLIP-GD [9] and StyleMC [11]. As shown in Fig. 11, the results produced by TediGAN are highly entangled and almost fail in editing attributes, such as “long hair” and “eyeglasses”. The results of StyleCLIP-LM and StyleMC are unstable and also entangled. Moreover, the results of StyleCLIP-GD have better disentanglement than StyleCLIP-LM as it is well-tunned for each case. However, the disadvantage of StyleCLIP-GD is the complex hyper-parameters tunning and long inference time as reported in Tab. 5. In comparison, DeltaEdit-G yields the most impressive and disentangled results almost in all cases.
@@ 5.2.3 Quantitative Comparison
Table 1: Quantitative comparison results of GAN-based methods.
@@
    StyleCLIP-GD ($\rho=2$)     $12.06$     $22.31$     $0.86$
    StyleCLIP-GD ($\rho=3$)     $16.85$     $19.50$     $0.77$
    StyleCLIP-GD ($\rho=4$)     $22.85$     $17.61$     $0.68$
@@
    DeltaEdit-G ($\mathcal{S}$ space)     $18.88$     $13.49$     $0.61$
    DeltaEdit-G ($\mathcal{W+}$ space)     $21.63$     $12.83$     $0.55$
    DeltaEdit-G (Straightforward)     $44.62$     $11.21$     $0.29$
In Tab. 1, we present objective measurements of FID, PSNR, and IDS (identity similarity before and after editing by Arcface [67]) for GAN-based methods comparison. All results are the average on ten given texts. Compared with the state-of-the-art GAN-based methods, our DeltaEdit-G achieves the best performance on all metrics.
We also compare DeltaEdit-G with the straightforward solution to text-free training in Tab. 1. For fairness, the straightforward solution is also trained in the $\mathcal{S}$ space and supervised by the same loss functions as the final solution. Compared with the straightforward solution, the final solution (ours) can largely improve the inference performance on all metrics, benefiting from the well-aligned CLIP Delta image-text space.
Table 2: User preference study on editing accuracy (Acc.) and visual realism (Real.)
@@
Finally, we conduct human subjective evaluations upon the editing accuracy (Acc) and visual realism (Real). We compare DeltaEdit-G with the strong GAN-based baseline, StyleCLIP-GD, under different hyper-parameter $\rho$ settings. In total, 20 evaluation rounds are performed and 40 participants are invited. At each round, we present results of randomly sampled editing text to each participant. Participants were asked to choose the best text-guided image editing output considering Acc and to select the output images (not limited to 1 image) which are visually realistic (Real). The results are listed in Tab. 2, showing the superiority of our model.
@@ 5.3.1 Qualitative Results
Table 3: Quantitative results of image reconstruction using SSIM, LPIPS, and MSE metrics.
@@
Real Image Reconstruction. The quality of image reconstruction can reflect the accuracy of a generative model in encoding-decoding real images. As shown in Tab. 3, we evaluate image reconstruction accuracy of various methods, including: (1) GAN inversion methods (including pretrained StyleGAN with $\mathcal{W}$ or $\mathcal{W+}$ sapce [18] as well as VQ-GAN [70]); (2) VAE-based methods (such as VQ-VAE2 [71]); (3) diffusion-based methods (including DDIM [53], DiffAE [54], and our proposed diffusion model). All models are trained on the FFHQ [17] dataset and tested on the CelebA-HQ [60] 30K dataset. All diffusion-based methods use T=100 for decoding. Our diffusion method and the diffusion-based DiffAE approach demonstrate superior performance, surpassing previous GAN-based and VAE-based methods by a large margin, thus confirming their near-perfect reconstruction capabilities.
@@ 5.3.2 Qualitative Comparison
Table 4: Quantitative comparison results of DeltaEdit-D with diffusion-based methods.
@@ 5.3.3 Quantitative Comparison
In Tab. 4, we present objective measurements of FID, PSNR, and IDS (identity similarity before and after editing by Arcface [67]) for diffusion-based methods comparison. The results are the average on ten given texts. The quantitative results have the best performance across multiple metrics. These results are consistent with the qualitative results that our DeltaEdit-D method exhibits the best disentanglement and identity preservation compared to diffusion-based models.
@@ 5.4 Comparison between DeltaEdit-G and DeltaEdit-D
Table 5: Time efficiency comparison of our method and other state-of-the-art methods. * means that the additional time for manually hyper-parameters tuning during the inference is excluded, which is typically 8-9 seconds for each case.
@@
5.5 Model Analysis
5.5.1 Text-prompt Setting
@@ 5.5.2 Efficiency Analysis
To validate the efficiency and flexibility of the proposed method, we compare the computation time with TediGAN, StyleMC and three StyleCLIP approaches in Tab. 5. Specifically, TediGAN first requires 12+ hours to encode images and texts into a common space and then trains the encoding module. StyleCLIP-OP manipulates images with an optimization process, which requires several tens of seconds. StyleCLIP-LM is fast during inference, but needs 10-12 hours to train a mapper network for per text prompt. StyleCLIP-GD requires about 4 hours to pre-compute the global edit directions. However, since the pre-computed directions are rough and cannot be directly applied to control the manipulation, it requires additional manual tuning for different text prompts, which typically takes 8-9 seconds for each case. Differently, StyleMC is fast during training. However, for N text prompts, StyleMC still needs N times longer to train the corresponding model. On the contrary, our DeltaEdit-G method avoids the labor time during training and inference. Once trained, our Delta Mapper is universal and can directly manipulate images with new text prompts efficiently. Meanwhile, our Delta Mapper only needs 2-3 hours for training, since the training is directly conducted on the latent space without generating images in each iteration. However, it is worth noting that DeltaEdit-D requires additional training time to learn the style-based diffusion model, resulting in a longer training phase compared to DeltaEdit-G. Additionally, since DeltaEdit-D relies on iterative denoising via the diffusion model to generate edited results, its inference time is also longer. Nevertheless, this progressive denoising process enables DeltaEdit-D to achieve higher-quality image edits than DeltaEdit-G, as demonstrated in Fig. 16.
@@ 5.5.3 Choice of the Editing Space
To find an appropriate space for the proposed method, we conduct experiments by performing editing in $\mathcal{W+}$ space and $\mathcal{S}$ space, validating the results on both GAN and diffusion models.. The quantitative and qualitative results in Tab. 1 and Fig. 18 show that $\mathcal{S}$ results can achieve better visual quality and identity preservation than $\mathcal{W+}$. For instance, in Fig. 18, when considering the attribute “bangs”, $\mathcal{W+}$ results succeed in adding bangs to given faces. However, the attribute-irrelevant contents are heavily changed, such as poses, skin color, and background. In contrast, $\mathcal{S}$ results have more disentanglement than $\mathcal{W+}$, e.g., the poses, facial identities are well preserved. Thus, we implement our DeltaEdit in $\mathcal{S}$ space, and introduce $R_{s}$ in the disentanglement part of Sec. 3.2 to further improve the disentanglement performance.
@@ 5.5.4 Effectiveness of Relevance Matrix $R_{s}$
For further improving the disentanglement, we introduce relevance matrix $R_{s}$ to limit some irrelevant channels from changing. The fourth row in Fig. 18 shows that, with $R_{s}$, our method can successfully edit the desired attributes while preserving the text-irrelevant content unchanged. For example, for “bangs”, the generated results have added bangs while accurately preserving the background and facial pose.
@@ 6 Conclusion
In this paper, we have introduced DeltaEdit, a flexible framework for text-guided image editing. The core of our approach has been the discovery and utilization of the well-aligned CLIP DeltaSpace. During training, the framework has learned to map differences in CLIP visual features to corresponding directions in a generative model’s latent space. For inference, it then leverages differences in CLIP text features to predict these latent directions. This design has enabled training without reliance on costly image-text pair collections and ensures strong generalization to unseen text prompts for zero-shot editing. Extensive qualitative and quantitative experiments across different generative models, including both GANs and diffusion models, have demonstrated the superiority of DeltaEdit. Our method has shown compelling advantages in generating high-quality results, achieving efficiency in both training and inference, and maintaining robust generalization to arbitrary, unseen textual descriptions.